# On the Equivalence of Random Network Distillation, Deep Ensembles, and Bayesian Inference

### v1

@@ Abstract
Uncertainty quantification is central to safe and efficient deployments of deep learning models, yet many computationally practical methods lack lacking rigorous theoretical motivation. Random network distillation (RND) is a lightweight technique that measures novelty via prediction errors against a fixed random target. While empirically effective, it has remained unclear what uncertainties RND measures and how its estimates relate to other approaches, e.g. Bayesian inference or deep ensembles. This paper establishes these missing theoretical connections by analyzing RND within the neural tangent kernel framework in the limit of infinite network width. Our analysis reveals two central findings in this limit: (1) The uncertainty signal from RND—its squared self-predictive error—is equivalent to the predictive variance of a deep ensemble. (2) By constructing a specific RND target function, we show that the RND error distribution can be made to mirror the centered posterior predictive distribution of Bayesian inference with wide neural networks. Based on this equivalence, we moreover devise a posterior sampling algorithm that generates i.i.d. samples from an exact Bayesian posterior predictive distribution using this modified Bayesian RND model. Collectively, our findings provide a unified theoretical perspective that places RND within the principled frameworks of deep ensembles and Bayesian inference, and offer new avenues for efficient yet theoretically grounded uncertainty quantification methods.
@@ 1 Introduction
Quantifying predictive uncertainty remains a cornerstone of reliable machine learning and underpins applications from safe robotics to efficiently exploring agents and autonomous scientific discovery. Bayesian inference is widely regarded as a theoretical gold‐standard to this end [neal1996bayesian, goan2020bayesian] but its application to neural networks is typically intractable in practice, requiring approximations of simplified posteriors through variational inference [VI, kingmaAutoEncodingVariationalBayes2014, galDropoutBayesianApproximation2016a, blei2017variational] or complex sampling mechanisms through Markov chain Monte Carlo approaches [MCMC, chen2014stochastic, liu2016stein, garriga2021exact]. Deep ensembles [dietterich2000ensemble, lakshminarayananSimpleScalablePredictive2017] on the other hand maintain several independently initialized models to quantify predictive variance as uncertainty. Due to their simplicity and relative practical reliability, deep ensembles have become a widely established alternative to Bayesian approaches for uncertainty quantification in deep learning [abdarReviewUncertaintyQuantification2021]. However, both ensemble methods and approximate Bayesian methods typically incur substantial computational and memory costs, in particular for larger-scale models, motivating more efficient alternatives. RND [burda2018exploration] offers one such approach: by training a predictor network to mimic the outputs of a fixed, randomly initialized target network, RND produces a simple novelty or uncertainty signal via the squared prediction error. Random network distillation (RND) has seen empirical success in exploration, out-of-distribution detection, and continual learning [burda2018exploration, nikulin2023anti, matthews2024craftax], yet the theoretical understanding of the nature of its uncertainty estimates remains blurry. In particular, it is unclear how—or whether—the RND error relates to the principled uncertainties produced for example by Bayesian inference or deep ensembles.
In this paper, we establish these missing theoretical connections by analyzing random network distillation in the idealized setting of infinite network width. In particular, we establish a Gaussian process (GP) interpretation of the self-predictive RND errors in the limit of infinitely wide neural networks, drawing on NTK theory [jacotNeuralTangentKernel2020, leeWideNeuralNetworks2020]. Our three main contributions are:
@@ 2 Preliminaries
Moreover, lee2017deep show that both $f(x;\theta_{0})$ and $f(x;\theta_{\infty})$ are indeed GPs described by the neural network Gaussian process [NNGP, lee2017deep] $f(x;\theta_{0})\sim\mathcal{GP}(0,\kappa_{xx^{\prime}})$ and the converged GP defined in Theorem 2. {restatable}thmrndwidenndistribution[leeWideNeuralNetworks2020](Distribution of post-convergence neural network functions) Let $f(\mathcal{X}_{T};\theta_{\infty})$ be a NN as defined in Eq.(1), and let $\mathcal{X}_{T}$ be testpoints. For random initializations $\theta_{0}\sim\mathcal{N}(0,I)$, and in the limit $n\to\infty$, $f(\mathcal{X}_{T};\theta_{\infty})$ distributes as a Gaussian with mean and covariance given by
@@
where h.c. is the Hermitian conjugate of the preceding term. See also Appendix B.1.2 or leeWideNeuralNetworks2020. Note that the GP described in Theorem 2 represents the law by which an infinite ensemble of infinitely wide neural networks from i.i.d. initializations distributes after training on $(\mathcal{X},\mathcal{Y})$, but—as is—permits no Bayesian posterior interpretation, which is of the canonical form described in Eq. 4.
@@ 3 Equivalence of Random Network Distillation & Deep Ensembles
We then draw on the results of Theorem 2 to provide a probabilistic description of the self-predictive errors $\epsilon(x;\vartheta_{\infty},\psi_{0})$ of a converged RND model in the limit of infinite network width. {restatable}thmrnderrordist(Distribution of post-convergence RND errors) Under NTK parametrization, let $u(x;\vartheta_{\infty})$ be a converged prediction network in $t\to\infty$, with data $\mathcal{X}$ and fixed target network $g(\mathcal{X};\psi_{0})$. Let parameters $\vartheta_{0},\psi_{0}$ be drawn i.i.d. $\vartheta_{0},\psi_{0}\sim\mathcal{N}(0,I)$, with the resulting NNGP $u(x;\vartheta_{0})\sim\mathcal{GP}(0,\kappa^{u}(x,x^{\prime}))$ and $g(x;\psi_{0})\sim\mathcal{GP}(0,\kappa^{g}(x,x^{\prime}))$. The post-convergence RND error $\epsilon(\mathcal{X}_{T};\vartheta_{\infty},\psi_{0})$ is Gaussian with zero mean and covariance
@@
where $\kappa^{\epsilon}_{xx^{\prime}}=\kappa^{u}_{xx^{\prime}}+\kappa^{g}_{xx^{\prime}}$ is the covariance kernel of initialization errors $\epsilon(x;\vartheta_{0},\psi_{0})=u(x;\vartheta_{0})-g(x;\psi_{0})$. Proof sketch. The error function $u(x;\vartheta_{\infty})-g(x;\psi_{0})$ is a sum of the random post-convergence function $u(x;\vartheta_{\infty})$ and the fixed random target function $g(x;\psi_{0})$. The latter $g(x;\psi_{0})$ is known to follow the NNGP. By the linearity of NTK learning dynamics, the online function $u(x;\vartheta_{\infty})$ is an affine transformation of its initialization $u(x;\vartheta_{0})$, which itself follows the NNGP. Moreover, this affine transformation is independent of $g$ or $\psi_{0}$, such that the error $\epsilon(x;\vartheta_{\infty},\psi_{0})$ is a sum of two independent GPs and therefore a GP itself. The resulting GP has zero-mean and covariance with an altered prior NNGP kernel $\kappa^{\epsilon}(x,x^{\prime})$ composed of the online prior kernel $\kappa^{u}_{xx^{\prime}}$ and the target prior kernel $\kappa^{g}_{xx^{\prime}}$. See also Appendix B.1.3.
{restatable}
crlrnderrorexp(Equivalence in expectation between RND errors and ensemble variance) Under the conditions of Theorem 3, let $\epsilon(x;\vartheta_{\infty},\psi_{0})$ be the error function of a converged RND network with data $\mathcal{X}$. Moreover, for a regression problem on $\mathcal{X}$ for some labels $\mathcal{Y}$, let $\mathbb{V}[f(x;\theta_{\infty})]$ denote the variance of converged NN functions random initializations. Furthermore, suppose an architectural equivalence between $f$, $u$, and $g$ and i.i.d. parameter initialization $\theta_{0},\vartheta_{0},\psi_{0}\sim\mathcal{N}(0,I)$. The expected norm of the RND error $\epsilon^{2}(x;\vartheta_{\infty},\psi_{0})$ then coincides with the ensemble variance
@@
Proof sketch. Corollary 3 follows straighforwardly from Theorem 3 by using $\kappa^{u}(x,x^{\prime})=\kappa^{g}(x,x^{\prime})$. Taking the trace of the covariance matrix and dividing by $2$, we recover the predictive ensemble variance $\mathbb{V}_{\theta_{0}}[f(x;\theta_{\infty})]$.
Theorem 3 and Corollary 3 formally show that, for an architectural equivalence between ensemble, predictor and target network, the expected RND errors directly quantify the predictive variance of the corresponding infinite ensemble model described by Theorem 2. To the best of our knowledge, it is the first formal analysis of random network distillation in the NTK regime and reveals a first theoretical motivation for the popular algorithm: in the idealized infinite-width setting, expected RND errors exactly quantify the variance of deep ensembles for any input $x$.
@@ 3.1 Multi-Headed Random Network Distillation
{restatable}
prprndindependenceprp(Independence of NN functions) Under NTK parametrization and in the limit $n\to\infty$, the random functions $f_{i}(x;\theta_{t})$ of a NN with $K$ output dimensions and shared hidden layers are mutually independent with covariance
@@
Notably, this decoupling holds despite the shared hidden layers and is an artifact of the learning dynamics exhibited in the infinite width limit and the NTK regime. In the absence of feature learning, output functions become statistically independent despite sharing a network body. By virtue of this independence property, a translation of the earlier obtained single-function results on RND error distributions (Theorem 3 and Corollary 3) to the multi-headed setting is straightforward. Our next result thus establishes an equivalence between the errors of the multi-headed RND algorithm, a widely used architecture in practice, and the variance of a finite-sized deep ensemble.
{restatable}
thmrndequivalence(Distributional equivalence between multi-headed RND and finite deep ensembles) Under the conditions of Theorem 3, let $u_{i}(x;\vartheta_{\infty}),g_{i}(x;\psi_{0})$ be the $i$-th output of predictor and target networks respectively with $K$ output dimensions. Denote their sample mean RND error $\bar{\epsilon}^{2}(x;\vartheta_{\infty},\psi_{0})=\frac{1}{K}\sum_{i=1}^{K}\epsilon_{i}^{2}(x;\vartheta_{\infty},\psi_{0})$. Moreover, let $\{f(x;\theta^{i}_{\infty})\}_{i=1}^{K+1}$ be an ensemble of $K+1$ NNs from i.i.d. initial draws $\theta_{0}$. Denote its sample variance $\bar{\sigma}_{f}^{2}(x;\theta_{\infty}^{i\dots K+1})=\frac{1}{K}\sum_{i=1}^{K+1}(f(x;\theta^{i}_{\infty})-\frac{1}{K+1}\sum_{j=1}^{K+1}f(x;\theta^{j}_{\infty}))^{2}$. The sample mean RND error and sample ensemble variance distribute to the same law
@@
where $\overset{D}{=}$ indicates an equality in distribution, namely by a scaled Chi-squared distribution $\bar{\sigma}_{f}^{2}(x;\theta_{\infty}^{i\dots K+1})\sim\frac{\Sigma^{f}_{xx}(\theta_{\infty})}{K}\chi^{2}(K)$ with scale $\Sigma^{f}_{xx}(\theta_{\infty})$ given by the analytical variance as given in Theorem 2.
@@
By Proposition 3.1, the function heads $\{u_{i}(x;\vartheta_{\infty})\}_{i=1}^{K}$ are $K$ independent predictors, each trained to match their independent targets $g_{i}(x;\psi_{0})$. Thus, the errors $\{\epsilon_{i}(x;\vartheta_{\infty},\psi_{0})\}_{i=1}^{K}$ are i.i.d. samples from the error distribution outlined in Proposition 3. In particular, $\bar{\epsilon}^{2}$ is the empirical mean of i.i.d. samples from a Gaussian which is known to be Chi-squared distributed. Similarly, we have that the ensemble $\{f(x;\theta^{i}_{\infty})\}_{i=1}^{K+1}$ are $K+1$ i.i.d. samples from the GP defined in Theorem 2, again yielding the known Chi-squared distribution for its sample variance $\bar{\sigma}_{f}^{2}(x;\theta_{\infty}^{i\dots K+1})$. See Appendix B.1.5.
@@ 4 Equivalence of Random Network Distillation & Bayesian Posteriors
When contrasting this identity with the GP governing the distribution of converged NN functions of Theorem 3, one observes a disparity in the structure of the covariance functions. While Theorem 2 and Theorem 3, too, specify GPs, they do not permit an interpretation as a Bayesian posterior predictive distribution [leeWideNeuralNetworks2020] due to the presence of two (in general) distinct kernel functions, namely the NNGP kernel $\kappa$ and the NTK $\Theta$. However, inspection of Theorem (3) and Eq. (10) suggests a path: if the prior kernel components within $\Sigma^{\epsilon}_{xx^{\prime}}(\vartheta_{\infty},\psi_{0})$, namely $\kappa_{xx^{\prime}}^{\epsilon}$, are be aligned with the dynamics kernel $\Theta_{xx^{\prime}}$ (i.e., if $\kappa^{\epsilon}\propto\Theta$), then the resulting covariance structure simplifies to the desired Bayesian posterior form of
@@
In the random network distillation algorithm, the prior kernel $\kappa^{\epsilon^{b}}_{xx^{\prime}}$ of initialization errors $\epsilon^{b}(x;\vartheta_{0},\vartheta_{0},\psi_{0})=u(x;\vartheta_{0})-\tilde{g}(x;\vartheta_{0},\psi_{0})$ is given by the sum of the online prior kernel and the target prior kernel $\kappa^{\epsilon^{b}}_{xx^{\prime}}=\kappa^{u}_{xx^{\prime}}+\kappa^{\tilde{g}}_{xx^{\prime}}$ (cf. Theorem 3), provided that $u$ and $\tilde{g}$ follow independent GPs. To obtain an error prior kernel that aligns with the NTK such that $\kappa^{\epsilon^{b}}_{xx^{\prime}}=\Theta_{xx^{\prime}}$, one may thus construct the target prior such that it satisfies $\kappa^{\tilde{g}}_{xx^{\prime}}=\Theta_{xx^{\prime}}-\kappa^{u}_{xx^{\prime}}$. To this end, a closer inspection of the relation between the NNGP kernel $\kappa^{u}_{xx^{\prime}}$ and the NTK $\Theta_{xx^{\prime}}$ is instructive. For this purpose, we will view the online network $u(x;\vartheta_{0})$ as a random feature model with its forward computation path as described in Eq. 1. Let in this scenario $x^{L}(x)$ denote the output vector, or the post-activations, before the final linear layer and denote the last-layer parameters at initialization $t=0$ as $(w^{L},b^{L})$. We can write the NN output at initialization $u(x;\vartheta_{0})$ as
@@
{restatable}
prprndntkgpinit (Bayesian RND target function) Under the conditions of Theorem 3, let $u(x;\vartheta_{0})$ and $g(x;\psi_{0})$ be neural networks of $L$ layers with parameters $\vartheta_{0},\psi_{0}\sim\mathcal{N}(0,I)$ i.i.d. Moreover, let $\psi_{0}^{L}=\{w^{L},b^{L}\}$ denote the last-layer parameters of $\psi_{0}$ and $\psi_{0}^{\leq L-1}$ the parameters of all preceding layers. Suppose the target function $\tilde{g}(x;\vartheta_{0},\psi_{0})$ is given by
@@
While the specific form of the kernel $\Theta_{xx^{\prime}}^{\leq L-1}=\Theta_{xx^{\prime}}-\Theta_{xx^{\prime}}^{L}$ seems unusual as a standalone prior, it is crucially important in shaping the final error distribution. This is because with the altered “Bayesian” target function $\tilde{g}(x;\vartheta_{0},\psi_{0})$ we can shape the covariance structure of errors at initialization by satisfying Eq. 14, appealing to Theorem (3). With the engineered target function $\tilde{g}(x;\vartheta_{0},\psi_{0})$, the learning dynamics of an RND model where the predictor network $u(x;\vartheta_{t})$ learns to mimic $\tilde{g}(\mathcal{X};\vartheta_{0},\psi_{0})$ can be shaped in the desired way. Our central statement is that the distribution of the error between the converged predictor $u(x;\vartheta_{\infty})$ and the target function $\tilde{g}(x;\vartheta_{0},\psi_{0})$ will then no longer reflect the variance of deep ensembles trained with gradient descent, but will instead directly exhibit the statistics of a Bayesian posterior predictive distribution derived from the NTK-GP prior. Theorem 4 formalizes this result.
{restatable}
thmrndbayesrnderrordist(Distribution of Bayesian RND errors) Under the conditions of Theorem 3, let $u(x;\vartheta_{\infty})$ be a converged predictor network trained on data $\mathcal{X}$ with labels from the fixed target function $\tilde{g}(\mathcal{X};\vartheta_{0},\psi_{0})$ as defined in Proposition 4. Let parameters $\vartheta_{0},\psi_{0}$ be drawn i.i.d. $\vartheta_{0},\psi_{0}\sim\mathcal{N}(0,I)$. The convergenced Bayesian RND error $\epsilon^{b}(\mathcal{X}_{T};\vartheta_{\infty},\vartheta_{0},\psi_{0})=u(\mathcal{X}_{T};\vartheta_{\infty})-\tilde{g}(\mathcal{X}_{T};\vartheta_{0},\psi_{0})$ on a test set $\mathcal{X}_{T}$ is Gaussian with zero mean and covariance

### v2

# Deep Ensembles, and Bayesian Inference

@@ Abstract
Uncertainty quantification is central to safe and efficient deployments of deep learning models, yet many computationally practical methods lack lacking rigorous theoretical motivation. Random network distillation (RND) is a lightweight technique that measures novelty via prediction errors against a fixed random target. While empirically effective, it has remained unclear what uncertainties RND measures and how its estimates relate to other approaches, e.g., Bayesian inference or deep ensembles. We establish these missing theoretical connections by analyzing RND within the neural tangent kernel framework in the limit of infinite network width. Our analysis reveals two central findings in this limit: (1) The uncertainty signal from RND—its squared self-predictive error—is equivalent to the predictive variance of a deep ensemble. (2) By constructing a specific RND target function, we show that the RND error distribution can be made to mirror the centered posterior predictive distribution of Bayesian inference with wide neural networks. Based on this equivalence, we moreover devise a posterior sampling algorithm that generates i.i.d. samples from an exact Bayesian posterior predictive distribution using this modified Bayesian RND model. Collectively, our findings provide a unified theoretical perspective that places RND within the principled frameworks of deep ensembles and Bayesian inference, and offer new avenues for efficient yet theoretically grounded uncertainty quantification methods.
@@ 1 Introduction
Quantifying predictive uncertainty remains a cornerstone of reliable machine learning and underpins applications from safe robotics to efficiently exploring agents and autonomous scientific discovery. Bayesian inference is widely regarded as a theoretical gold‐standard to this end [neal1996bayesian, goan2020bayesian] but its application to neural networks is typically intractable in practice, requiring approximations of simplified posteriors through variational inference [VI, kingmaAutoEncodingVariationalBayes2014, galDropoutBayesianApproximation2016a, blei2017variational] or complex sampling mechanisms through Markov chain Monte Carlo approaches [MCMC, chen2014stochastic, liu2016stein, garriga2021exact]. Deep ensembles [dietterich2000ensemble, lakshminarayananSimpleScalablePredictive2017] on the other hand maintain several independently initialized models to quantify predictive variance as uncertainty. Due to their simplicity and relative practical reliability, deep ensembles have become a widely established alternative to Bayesian approaches for uncertainty quantification in deep learning [abdarReviewUncertaintyQuantification2021].
However, both ensemble methods and approximate Bayesian methods typically incur substantial computational and memory costs, in particular for larger-scale models, motivating more efficient alternatives. RND [burda2018exploration] offers one such approach: by training a predictor network to mimic the outputs of a fixed, randomly initialized target network, RND produces a simple novelty or uncertainty signal via the squared prediction error. Random network distillation (RND) has seen empirical success in exploration, out-of-distribution detection, and continual learning [burda2018exploration, nikulin2023anti, matthews2024craftax], yet the theoretical understanding of the nature of its uncertainty estimates remains blurry. In particular, it is unclear how—or whether—the RND error relates to the principled uncertainties produced for example by Bayesian inference or deep ensembles.
In this paper, we establish these missing theoretical connections by analyzing random network distillation in the idealized setting of infinite network width. In particular, we establish a Gaussian process (GP) interpretation of the self-predictive RND errors in the limit of infinitely wide neural networks, drawing on Neural Tangent Kernel (NTK) theory [jacotNeuralTangentKernel2020, leeWideNeuralNetworks2020]. Our three main contributions are:
@@ 2 Preliminaries
Moreover, lee2017deep show that both $f(x;\theta_{0})$ and $f(x;\theta_{\infty})$ are indeed GPs described by the neural network Gaussian process [NNGP, lee2017deep] $f(x;\theta_{0})\sim\mathcal{GP}(0,\kappa_{xx^{\prime}})$ and the converged GP defined in Theorem 2.1.
Theorem 2.1.
[leeWideNeuralNetworks2020](Distribution of post-convergence neural network functions) Let $f(\mathcal{X}_{T};\theta_{\infty})$ be a NN as defined in Eq.(1), and let $\mathcal{X}_{T}$ be testpoints. For random initializations $\theta_{0}\sim\mathcal{N}(0,I)$, and in the limit $n\to\infty$, $f(\mathcal{X}_{T};\theta_{\infty})$ distributes as a Gaussian with mean and covariance given by
@@
where h.c. is the Hermitian conjugate of the preceding term.
See also Appendix B.1.2 or leeWideNeuralNetworks2020. Note that the GP described in Theorem 2.1 represents the law by which an infinite ensemble of infinitely wide neural networks from i.i.d. initializations distributes after training on $(\mathcal{X},\mathcal{Y})$, but—as is—permits no Bayesian posterior interpretation, which is of the canonical form described in Eq. 4.
@@ 3 Equivalence of Random Network Distillation & Deep Ensembles
We then draw on the results of Theorem 2.1 to provide a probabilistic description of the self-predictive errors $\epsilon(x;\vartheta_{\infty},\psi_{0})$ of a converged RND model in the limit of infinite network width.
Theorem 3.1.
(Distribution of post-convergence RND errors) Under NTK parametrization, let $u(x;\vartheta_{\infty})$ be a converged prediction network in $t\to\infty$, with data $\mathcal{X}$ and fixed target network $g(\mathcal{X};\psi_{0})$. Let parameters $\vartheta_{0},\psi_{0}$ be drawn i.i.d. $\vartheta_{0},\psi_{0}\sim\mathcal{N}(0,I)$, with the resulting NNGP $u(x;\vartheta_{0})\sim\mathcal{GP}(0,\kappa^{u}(x,x^{\prime}))$ and $g(x;\psi_{0})\sim\mathcal{GP}(0,\kappa^{g}(x,x^{\prime}))$. The post-convergence RND error $\epsilon(\mathcal{X}_{T};\vartheta_{\infty},\psi_{0})$ is Gaussian with zero mean and covariance
@@
where $\kappa^{\epsilon}_{xx^{\prime}}=\kappa^{u}_{xx^{\prime}}+\kappa^{g}_{xx^{\prime}}$ is the covariance kernel of initialization errors $\epsilon(x;\vartheta_{0},\psi_{0})=u(x;\vartheta_{0})-g(x;\psi_{0})$.
Proof sketch. The error function $u(x;\vartheta_{\infty})-g(x;\psi_{0})$ is a sum of the random post-convergence function $u(x;\vartheta_{\infty})$ and the fixed random target function $g(x;\psi_{0})$. The latter $g(x;\psi_{0})$ is known to follow the NNGP. By the linearity of NTK learning dynamics, the online function $u(x;\vartheta_{\infty})$ is an affine transformation of its initialization $u(x;\vartheta_{0})$, which itself follows the NNGP. Moreover, this affine transformation is independent of $g$ or $\psi_{0}$, such that the error $\epsilon(x;\vartheta_{\infty},\psi_{0})$ is a sum of two independent GPs and therefore a GP itself. The resulting GP has zero-mean and covariance with an altered prior NNGP kernel $\kappa^{\epsilon}(x,x^{\prime})$ composed of the online prior kernel $\kappa^{u}_{xx^{\prime}}$ and the target prior kernel $\kappa^{g}_{xx^{\prime}}$. See also Appendix B.1.3.
Corollary 3.2.
(Equivalence in expectation between RND errors and ensemble variance) Under the conditions of Theorem 3.1, let $\epsilon(x;\vartheta_{\infty},\psi_{0})$ be the error function of a converged RND network with data $\mathcal{X}$. Moreover, for a regression problem on $\mathcal{X}$ for some labels $\mathcal{Y}$, let $\mathbb{V}[f(x;\theta_{\infty})]$ denote the variance of converged NN functions random initializations. Furthermore, suppose an architectural equivalence between $f$, $u$, and $g$ and i.i.d. parameter initialization $\theta_{0},\vartheta_{0},\psi_{0}\sim\mathcal{N}(0,I)$. The expected norm of the RND error $\epsilon^{2}(x;\vartheta_{\infty},\psi_{0})$ then coincides with the ensemble variance
@@
Proof sketch. Corollary 3.2 follows straighforwardly from Theorem 3.1 by using $\kappa^{u}(x,x^{\prime})=\kappa^{g}(x,x^{\prime})$. Taking the trace of the covariance matrix and dividing by $2$, we recover the predictive ensemble variance $\mathbb{V}_{\theta_{0}}[f(x;\theta_{\infty})]$.
Theorem 3.1 and Corollary 3.2 formally show that, for an architectural equivalence between ensemble, predictor and target network, the expected RND errors directly quantify the predictive variance of the corresponding infinite ensemble model described by Theorem 2.1. To the best of our knowledge, it is the first formal analysis of random network distillation in the NTK regime and reveals a first theoretical motivation for the popular algorithm: in the idealized infinite-width setting, expected RND errors exactly quantify the variance of deep ensembles for any input $x$.
@@ 3.1 Multi-Headed Random Network Distillation
Proposition 3.3.
(Independence of NN functions) Under NTK parametrization and in the limit $n\to\infty$, the random functions $f_{i}(x;\theta_{t})$ of a NN with $K$ output dimensions and shared hidden layers are mutually independent with covariance
@@
Notably, this decoupling holds despite the shared hidden layers and is an artifact of the learning dynamics exhibited in the infinite width limit and the NTK regime. In the absence of feature learning, output functions become statistically independent despite sharing a network body. By virtue of this independence property, a translation of the earlier obtained single-function results on RND error distributions (Theorem 3.1 and Corollary 3.2) to the multi-headed setting is straightforward. Our next result thus establishes an equivalence between the errors of the multi-headed RND algorithm, a widely used architecture in practice, and the variance of a finite-sized deep ensemble.
Theorem 3.4.
(Distributional equivalence between multi-headed RND and finite deep ensembles) Under the conditions of Theorem 3.1, let $u_{i}(x;\vartheta_{\infty}),g_{i}(x;\psi_{0})$ be the $i$-th output of predictor and target networks respectively with $K$ output dimensions. Denote their sample mean RND error $\bar{\epsilon}^{2}(x;\vartheta_{\infty},\psi_{0})=\frac{1}{K}\sum_{i=1}^{K}\epsilon_{i}^{2}(x;\vartheta_{\infty},\psi_{0})$. Moreover, let $\{f(x;\theta^{i}_{\infty})\}_{i=1}^{K+1}$ be an ensemble of $K+1$ NNs from i.i.d. initial draws $\theta_{0}$. Denote its sample variance $\bar{\sigma}_{f}^{2}(x;\theta_{\infty}^{i\dots K+1})=\frac{1}{K}\sum_{i=1}^{K+1}(f(x;\theta^{i}_{\infty})-\frac{1}{K+1}\sum_{j=1}^{K+1}f(x;\theta^{j}_{\infty}))^{2}$. The sample mean RND error and sample ensemble variance distribute to the same law
@@
where $\overset{D}{=}$ indicates an equality in distribution, namely by a scaled Chi-squared distribution $\bar{\sigma}_{f}^{2}(x;\theta_{\infty}^{i\dots K+1})\sim\frac{\Sigma^{f}_{xx}(\theta_{\infty})}{K}\chi^{2}(K)$ with scale $\Sigma^{f}_{xx}(\theta_{\infty})$ given by the analytical variance as given in Theorem 2.1.
@@
By Proposition 3.3, the function heads $\{u_{i}(x;\vartheta_{\infty})\}_{i=1}^{K}$ are $K$ independent predictors, each trained to match their independent targets $g_{i}(x;\psi_{0})$. Thus, the errors $\{\epsilon_{i}(x;\vartheta_{\infty},\psi_{0})\}_{i=1}^{K}$ are i.i.d. samples from the error distribution outlined in Proposition 3.2. In particular, $\bar{\epsilon}^{2}$ is the empirical mean of i.i.d. samples from a Gaussian which is known to be Chi-squared distributed. Similarly, we have that the ensemble $\{f(x;\theta^{i}_{\infty})\}_{i=1}^{K+1}$ are $K+1$ i.i.d. samples from the GP defined in Theorem 2.1, again yielding the known Chi-squared distribution for its sample variance $\bar{\sigma}_{f}^{2}(x;\theta_{\infty}^{i\dots K+1})$. See Appendix B.1.5.
@@ 4 Equivalence of Random Network Distillation & Bayesian Posteriors
When contrasting this identity with the GP governing the distribution of converged NN functions of Theorem 3.1, one observes a disparity in the structure of the covariance functions. While Theorem 2.1 and Theorem 3.1, too, specify GPs, they do not permit an interpretation as a Bayesian posterior predictive distribution [leeWideNeuralNetworks2020] due to the presence of two (in general) distinct kernel functions, namely the NNGP kernel $\kappa$ and the NTK $\Theta$. However, inspection of Theorem (3.1) and Eq. (10) suggests a path: if the prior kernel components within $\Sigma^{\epsilon}_{xx^{\prime}}(\vartheta_{\infty},\psi_{0})$, namely $\kappa_{xx^{\prime}}^{\epsilon}$, are be aligned with the dynamics kernel $\Theta_{xx^{\prime}}$ (i.e., if $\kappa^{\epsilon}\propto\Theta$), then the resulting covariance structure simplifies to the desired Bayesian posterior form of
@@
In the random network distillation algorithm, the prior kernel $\kappa^{\epsilon^{b}}_{xx^{\prime}}$ of initialization errors $\epsilon^{b}(x;\vartheta_{0},\vartheta_{0},\psi_{0})=u(x;\vartheta_{0})-\tilde{g}(x;\vartheta_{0},\psi_{0})$ is given by the sum of the online prior kernel and the target prior kernel $\kappa^{\epsilon^{b}}_{xx^{\prime}}=\kappa^{u}_{xx^{\prime}}+\kappa^{\tilde{g}}_{xx^{\prime}}$ (cf. Theorem 3.1), provided that $u$ and $\tilde{g}$ follow independent GPs. To obtain an error prior kernel that aligns with the NTK such that $\kappa^{\epsilon^{b}}_{xx^{\prime}}=\Theta_{xx^{\prime}}$, one may thus construct the target prior such that it satisfies $\kappa^{\tilde{g}}_{xx^{\prime}}=\Theta_{xx^{\prime}}-\kappa^{u}_{xx^{\prime}}$. To this end, a closer inspection of the relation between the NNGP kernel $\kappa^{u}_{xx^{\prime}}$ and the NTK $\Theta_{xx^{\prime}}$ is instructive. For this purpose, we will view the online network $u(x;\vartheta_{0})$ as a random feature model with its forward computation path as described in Eq. 1. Let in this scenario $x^{L}(x)$ denote the output vector, or the post-activations, before the final linear layer and denote the last-layer parameters at initialization $t=0$ as $(w^{L},b^{L})$. We can write the NN output at initialization $u(x;\vartheta_{0})$ as
@@
Proposition 4.1.
(Bayesian RND target function) Under the conditions of Theorem 3.1, let $u(x;\vartheta_{0})$ and $g(x;\psi_{0})$ be neural networks of $L$ layers with parameters $\vartheta_{0},\psi_{0}\sim\mathcal{N}(0,I)$ i.i.d. Moreover, let $\psi_{0}^{L}=\{w^{L},b^{L}\}$ denote the last-layer parameters of $\psi_{0}$ and $\psi_{0}^{\leq L-1}$ the parameters of all preceding layers. Suppose the target function $\tilde{g}(x;\vartheta_{0},\psi_{0})$ is given by
@@
While the specific form of the kernel $\Theta_{xx^{\prime}}^{\leq L-1}=\Theta_{xx^{\prime}}-\Theta_{xx^{\prime}}^{L}$ seems unusual as a standalone prior, it is crucially important in shaping the final error distribution. This is because with the altered “Bayesian” target function $\tilde{g}(x;\vartheta_{0},\psi_{0})$ we can shape the covariance structure of errors at initialization by satisfying Eq. 14, appealing to Theorem (3.1). With the engineered target function $\tilde{g}(x;\vartheta_{0},\psi_{0})$, the learning dynamics of an RND model where the predictor network $u(x;\vartheta_{t})$ learns to mimic $\tilde{g}(\mathcal{X};\vartheta_{0},\psi_{0})$ can be shaped in the desired way. Our central statement is that the distribution of the error between the converged predictor $u(x;\vartheta_{\infty})$ and the target function $\tilde{g}(x;\vartheta_{0},\psi_{0})$ will then no longer reflect the variance of deep ensembles trained with gradient descent, but will instead directly exhibit the statistics of a Bayesian posterior predictive distribution derived from the NTK-GP prior. Theorem 4.2 formalizes this result.
Theorem 4.2.
(Distribution of Bayesian RND errors) Under the conditions of Theorem 3.1, let $u(x;\vartheta_{\infty})$ be a converged predictor network trained on data $\mathcal{X}$ with labels from the fixed target function $\tilde{g}(\mathcal{X};\vartheta_{0},\psi_{0})$ as defined in Proposition 4.1. Let parameters $\vartheta_{0},\psi_{0}$ be drawn i.i.d. $\vartheta_{0},\psi_{0}\sim\mathcal{N}(0,I)$. The convergenced Bayesian RND error $\epsilon^{b}(\mathcal{X}_{T};\vartheta_{\infty},\vartheta_{0},\psi_{0})=u(\mathcal{X}_{T};\vartheta_{\infty})-\tilde{g}(\mathcal{X}_{T};\vartheta_{0},\psi_{0})$ on a test set $\mathcal{X}_{T}$ is Gaussian with zero mean and covariance